Pablo Company calculates the cost for an equivalent unit of production using costing.

Data For June

Work in process inventory, June 1: 30,000 units

Direct materials: 100% complete 60,000

Conversation: 40% complete 25,000

Balance in work process 85,000

Units started during June 70,000

Units completed and transferred out 70,000

Work in process inventory, June 30 30,000

Direct materials: 100% complete

Conservation: 80% complete

Costs incurred during June

Direct materials 122,500

Conversation costs

Direct labor 119,720

Applied overhead 179,580

Total conversation costs 299,300

Compute the costs per equivalent unit for both the weighted- average and FIFO methods

.................................................Weighted Average Cost per EU...............FIFO cost per EU

Direct materials

Conversion

Total cost
Answer #1

Equivalent unit : Weighted average method

Physical Unit EUP-Direct material EUP-Conversion
Unit Completed and transferred out 70000 70000 70000
Unit ending work in process 30000 30000 30000*80% = 24000
Equivalent unit of production 100000 100000 94000

Equivalent unit : FIFO method

Physical Unit EUP-Direct material EUP-Conversion
Unit beginning work in process 30000 30000*0% = 0 30000*60% = 18000
Unit started and completed 40000 40000 40000
Unit ending work in process 30000 30000 30000*80% = 24000
Equivalent unit of production 100000 70000 82000

Calculate cost per equivalent unit

Weighted Average Cost per EU FIFO cost per EU
Direct material (60000+122500) = 182500/100000 = 1.825 122500/70000 = 1.75
Conversion (25000+299300) = 324300/94000 = 3.45 299300/82000 = 3.65
Total Cost $5.275 $5.40